OpenStudy (anonymous):

If f(x)=x^2+3x+5 what is f(3+h)

Directrix (directrix):

f(x)=x^2+3x+5 In f(3+h), the role of x is played by (3 + h). So, everywhere there is an x in f(x)=x^2+3x+5, replace that x by (3 + h). f(3 + h) = (3 + h)^2 + 3* (3 + h) + 5 Now, simplify: (3 + h)^2 + 3* (3 + h) + 5 @olamy